OpenStudy (anonymous):

If -nx^2 + tx + c = 0, what is x equal to?

OpenStudy (anonymous):

Answer choices attached

OpenStudy (freckles):

you can use the quadratic formula

OpenStudy (anonymous):

How would I do that?

OpenStudy (freckles):

\[\text{ if } ax^2+bx+c=0 \text{ then } x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Okay, I see Thank you :)

OpenStudy (anonymous):

\[x= -t \pm \sqrt{t^2-4(-n)(c)} \div 2(-n\] Would this be correct? @freckles

OpenStudy (freckles):

ok and you can clean it only a little bit really

OpenStudy (anonymous):

Thank you :)

OpenStudy (freckles):